Wholesome Foods sell dried apricots for $12.25 per pound and dried apples for $7.50 per pound. How much of each fruit should be used to make a 40-lb mixture that sells for $9.40 per pound?
Answer by stanbon(75887)

Wholesome Foods sell dried apricots for $12.25 per pound and dried apples for $7.50 per pound. How much of each fruit should be used to make a 40-lb mixture that sells for $9.40 per pound?
---------------
Equation:
value + value = value
12.25d + 7.50(40-d) = 9.40*40
-------
12.25d + 7.50*40-7.50d = 9.40*40
-----
4.75d = 1.9*40
-----
d = 16 lbs (amt. of dried apricots needed)
40-d = 24 lbs (amt of dried apples needed)
------------
Cheers,
Stan H.
----------